Use the information in the table below to answer:

CPI Index (Dec 2012 = 100)

Year July August September October

2014 104 106 107 110

2015 112 113 115 116

  1. The monthly inflation rate for October is 0.90 %. TRUE OR FALSE?

.2. The annual inflation rate for October 2015 is 5.45 %. TRUE OR FALSE?







1
Expert's answer
2021-09-02T12:35:54-0400

1.) "110-107 = 3"


"\frac{3}{107} ×100 = 2%" %


Therefore, it is False


2.) "116- 115 = 1"


"=\frac{ 1}{115 } × 100" "= 0.87" %


Therefore, it is False